# Differential pumping for kHz operation of a Laser Wakefield accelerator based on a continuously flowing Hydrogen gas jet

### v2

@@ Abstract
Laser-Wakefield Accelerators (LWFA) running at kHz repetition rates hold great potential for applications. They typically operate with low-energy, highly compressed laser pulses focused in high-pressure gas targets. Experiments have shown that the best-quality electron beams are achieved using Hydrogen gas targets. However, continuous operation with Hydrogen requires a dedicated pumping system. This work presents the design of a differential pumping system, enabling, for the first time, continuous operation of our kHz LWFA using a high-pressure Hydrogen gas jet. The system successfully maintained a pressure below $3\times 10^{-4}\,$mbar, even with a free-flowing gas jet operating at 140 bar backing pressure. Numerical fluid dynamics and optical simulations were used to guide and validate the system’s design.
@@ I Introduction
In LWFA, the laser first propagates through a vacuum before being focused into a plasma, where the acceleration process occurs. The plasma is generated as the laser ionizes the gas supplied by the target system. Various types of gas targets are used in LWFA. Each type is designed to balance the required plasma density while minimizing gas leakage, which increases the pressure in the vacuum chamber where the laser propagates. Here are the most commonly used target systems: (i) gas jets: a supersonic gas flow expands from a nozzle into the vacuum chamber Semushin and Malka (2001); Schmid and Veisz (2012), (ii) gas cells: the gas is enclosed in a cell, creating an interaction volume in which the laser propagates through two apertures at the cell’s entrance and exit Osterhoff et al. (2008). These apertures must be small enough to minimize gas leakage in the main chamber, but large enough to avoid ablation from the laser which would in turn negatively affect the laser performance. In order to mitigate the leaks, the gas cells operate at lower gas pressures, resulting in lower plasma densities compared to gas jets. Additionally, some gas cells utilize differential pumping, where the gas is removed from the cell simultaneously as it is introduced Kirchen et al. (2021); Drobniak et al. (2023). A last commonly used target system is the capillary discharge target, where the plasma is generated via a high-voltage discharge within a capillary filled with gas. There, the laser propagates through the guided mode of the plasma channel formed during plasma expansion and interaction with the capillary walls Leemans et al. (2006). This approach is relevant when long acceleration distances are needed (in the range of 111\,1cm), but the entrance capillary aperture can be damaged by high-intensity lasers. It is worth noticing that similar target systems are also being employed in the field of nuclear physics. Notably, some of their experiments make use of gas jets combined with differential pumping techniques, though these systems generally operate at much lower densities than those in the present study Ulbricht, Clausnitzer, and Graw (1972); Bittner, Kretschmer, and Schuster (1979); Treichel, Isenbügel, and Marquardt (1983); Shapira et al. (1985). Early-stage designs and implementations of such differential pumping systems have been explored in the frame of LWFA Bohlen et al. (2022).
Advances in laser technology, along with the development of the target systems, have led to tremendous progress in the laser-plasma accelerator performances. Today, LWFA driven by 100 TW to PW scale laser systems can produce electron beams with energies ranging from 100100100\,100MeV to GeV over a few centimeters Mangles et al. (2004); Geddes et al. (2004); Faure et al. (2004); Kim et al. (2017); Oubrerie et al. (2022a); Miao et al. (2022). To date, the highest energy ever reached is 8 GeV using plasma discharge technology Gonsalves et al. (2019).
@@
Regarding vacuum conditions, utilizing a nitrogen plasma is convenient: each $\mathrm{N_{2}}$ molecule provides 10 electrons, enabling a backing pressure of only 20 to 30 bar in the gas jets to achieve the required plasma density. Additionally, heavy gases like nitrogen are easier to pump than light gases Gaede (1915); Becker (1966). However, ionization effects can significantly distort the laser pulse in $\mathrm{N_{2}}$ as different levels are ionizes at different positions within the laser pulse, causing a decrease in accelerator performance. To counter this, utilizing a light gas such as hydrogen is necessary Monzac et al. (2024). So far, kHz-scale experiments using an $\mathrm{H_{2}}$ plasma have been limited to burst mode operation as pumping $\mathrm{H_{2}}$ poses significant challenges Salehi et al. (2017, 2021). When a free-flowing gas jet of $\mathrm{H_{2}}$ is used at high pressure, it results in considerable gas loading in the vacuum chamber where the experiment is conducted. This phenomenon was observed by Salehi et al. Salehi et al. (2021), who reported a rapid pressure increase in the main chamber, rising from $2.6\times 10^{-2}\,$mbar to $2.0\times 10^{-1}\,$mbar in just 1 second when using a free-flowing gas jet with a $\sim 150\,\text{\textmu}$ µm FWHM aperture and a backing pressure of $\sim 30\,$bar. They operated their accelerator in burst mode and avoided gas loading in their experiments.
@@ II Differential pumping with gas jets and high pressure light gases
The optimal plasma density and length for a laser-plasma accelerator depend on the physical properties of the laser pulse. In the case of a few mJ, few fs pulse, the plasma density should typically be $n_{e}\sim 10^{19}-10^{20}\,\mathrm{cm}^{-3}$ and the plasma length $L=100-200\,\text{\textmu}$ µm. To reach such plasma parameters, we use supersonic cylindrical gas nozzles and backing pressures ranging from a few tens of bar to about 150150150\,150bar. A cross-section diagram of the gas jet is shown in Figure 1b. In the following, we will refer to the gas nozzle parameters as follows: the diameter of the nozzle throat and nozzle opening are respectively written $D^{*}$ and $D$ and the corresponding throat and opening surfaces are $A^{*}$ and $A$, where we simply have $A=\pi D^{2}/4$. For example, a 60/180 gas nozzle has $D^{*}=60\,\text{\textmu}$ µm, $D=180\,\text{\textmu}$ µm, and provides the right plasma length and molecular density $n_{N_{2}}\sim 10^{19}\,\mathrm{cm}^{-3}$ when using $\mathrm{N_{2}}$ gas with pressure $P_{back}=30\,$bar backing pressure (see Figure 1a). The laser pulse is intense enough to ionize Nitrogen atoms to N5+, so that each $\mathrm{N_{2}}$ molecule releases 10 electrons, leading to the required plasma electron density of $n_{e}\sim 10^{20}\,\mathrm{cm}^{-3}$.
Figure 1: a) Electronic density $n_{e}$ of the plasma as a function of the backing pressure in the gas jet and the nature of the gas for a 60/180 supersonic nozzle. b) Cross section of the supersonic gas jet, $D^{*}$ and $D$ are respectively the diameter of the nozzle throat and nozzle opening.
@@ II.1 Differential pumping use in high repetition rate LWFA
The pumping system is designed so that the residual equilibrium pressure in the experimental chamber, $P_{eq}^{(1)}$ stays sufficiently low and does not affect laser propagation. We set this limit empirically to $P_{eq}^{(1)}<10^{-2}\,$mbar since above this value, turbo-molecular pumps become inoperative, overheat, and the pumping system diverges. The residual pressure at equilibrium $P_{eq}^{(1)}$ in the chamber can be estimated knowing the pumping speed $R_{eff}[\mathrm{l/s}]$ and the gas leak rate due to the gas jet $Q[\mathrm{mbar\cdot l/s}]$ by simply writing that:
@@
This allows to easily estimate the gas leak rate in our experiment. Typically, in Nitrogen, using a nozzle with $D^{*}=60\,\text{\textmu}$ µm and a backing pressure of $P_{back}=20\,$bar, we measure $P_{eq}^{(1)}=3\times 10^{-3}\,$mbar. Knowing that the pumping speed of our turbo-molecular pump is $R_{eff}=4000\,$l/s for Nitrogen, we obtain that the gas leak is $Q=12\,$mbar.l/s.
Let us now calculate what happens if one uses $\mathrm{H_{2}}$ instead of Nitrogen. We assume complete ionization: each $\mathrm{H_{2}}$ molecule now releases 2 electrons instead of 10 in $\mathrm{N_{2}}$. In addition, we use a 1D isentropic expansion model for compressible gas flows to extrapolate the molecular density in $\mathrm{H_{2}}$. The molecular density in the flow, $n_{mol}$, can be expressed according to the Boltzmann constant $k_{B}$, the Mach number $M_{Mach}$ and the initial pressure and temperature, $P_{0},\,T_{0}$ in the reservoir Zucker and Biblarz (2002):
@@
$Q_{H_{2}}=5Q_{N_{2}}\left(\frac{M_{N_{2}}}{M_{H_{2}}}\right)^{1/2}$ (3)
so that instead of $Q=12\,$mbar.l/s for 20 bar of $\mathrm{N_{2}}$, the leak rate would be $Q=224\,$mbar.l/s for 100 bar of $\mathrm{H_{2}}$. To compensate for that increase and keep the background pressure below $P_{eq}^{(1)}=3\times 10^{-3}\,$mbar, as when using $\mathrm{N_{2}}$, the pumping speed should be higher than $R_{eff}=7.4\times 10^{4}\,$l/s which is unrealistic as the largest turbo-molecular pumps on the market have maximum pumping speeds of 4000 l/s. Moreover, one should notice that $\mathrm{H_{2}}$ is more difficult to pump than $\mathrm{N_{2}}$. Overall, heavy gases are easier to pump than light gases; the compression ratio of the turbo-molecular pump goes as $K\propto\exp(\sqrt{M})$ with M the molar mass of the gas, and the pumping speed goes as $S\propto-\ln(K)$ Gaede (1915); Becker (1966), so the lighter the gas, the slower the pumping speed of the turbomolecular pump. This has been the bottleneck for using continuous flows of $\mathrm{He}$ or $\mathrm{H_{2}}$ in LWFA. For example, in our system, the pumping speed of our turbo-molecular pump drops down from 400040004000\,4000l/s in $\mathrm{N_{2}}$ to 300300300\,300l/s in $\mathrm{H_{2}}$ at $P=10^{-1}\,$mbar. In practice, this means that for $P_{0}=100\,$bar of $\mathrm{H_{2}}$, the equilibrium pressure would be around $P_{eq}^{(1)}\simeq 0.5\,$mbar, which is likely too high for the turbo-molecular pump to operate. Similar conclusions could be reached for $\mathrm{He}$.
Figure 2: Without differential pumping: equilibrium pressure in the main chamber versus backing pressure for a gas nozzle with throat diameter $D^{*}=100\,\text{\textmu}$ µm. The pumping system is a turbo-molecular pump 400040004000\,4000l/s.
This has been verified experimentally: we have measured the pressure in the vacuum chamber in which a supersonic nozzle lets out a continuous flow of gas. The vacuum chamber is pumped by a $R_{eff}=4000\,$l/s turbo-molecular pump connected to a primary pump with pumping speed 130130130\,130l/s. In Figure 2, we show the variation of the equilibrium pressure versus backing pressure for $\mathrm{N_{2}}$, and $\mathrm{He}$. The results show very clearly that for $\mathrm{He}$, the equilibrium pressure increases abruptly above $P_{back}=20\,$bar and the pumping system diverges above that, not being able to dispose of the high $\mathrm{He}$ flow. This clearly confirms that a specific solution should be implemented for pumping high-pressure ($P>100\,$bar) light gases.
@@ II.2 Design of the differential pumping system
An alternative solution is to use a differential pumping system: the gas jet is enclosed into a small vacuum chamber, the differential chamber, that is pumped by a powerful primary pump able to work efficiently at the mbar level with a pumping speed of 360360360\,360l/s. The idea is to pump out most of the gas from the small chamber while tolerating a higher equilibrium pressure at the mbar level, $P_{eq}^{(2)}\sim\,$mbar within the differential chamber subsystem. Then, by restricting the leakage from the differential chamber to the main chamber, the pressure can be kept below $P_{eq}^{(1)}<10^{-3}\,$mbar in the main chamber. In this case, the laser propagation is not affected by the residual gas, provided that the mbar region in the differential chamber is sufficiently short, e.g. millimeter scale. Figure 3 shows a conceptual schematic of the differential scheme that is implemented in the experiment. The laser enters and exits the chamber via two cones that terminate with 111\,1mm holes for restricting gas leakage. In addition, the “high” density region is limited to a $4$-mm gap in between the two cones to limit its impact on laser propagation.
@@
Interestingly, the experimental measurement of the gas leak rate $Q$ can be compared to a theoretical estimation using well-known fluid gas dynamics Zucker and Biblarz (2002). An estimation of the mass flow rate can be obtained using a 1D isentropic model of the gas flow in the nozzle. For a 1D stationary and turbulent flow, the mass flow rate $\dot{m}=\rho\nu A$ is constant throughout the flow; at the nozzle throat, it reduces to:
@@
where $\rho^{*},\,\nu^{*}$ and $A^{*}$ are respectively the mass density, fluid velocity, and the surface at the nozzle throat, $P_{0}$ and $T_{0}$ are the pressure and temperature in the reservoir. Note that here, subscript 0 refers to the reservoir, i.e. the stagnation volume where the fluid velocity is $\nu_{0}=0$. The leak rate in mbar.l/s can then be obtained from $Q\,=PdV/dt=k_{B}T\frac{dN}{dt}$ which can eventually be written as:
@@
[…]
@@
Using this model, it is straightforward to determine $P_{eq}^{(1)}$ and $P_{eq}^{(2)}$, respectively the equilibrium pressures in the experimental and differential chambers, that are reached as a function of the pumping speed $R_{eff}$ used for pumping the differential chamber. Figure 4 shows the results for a backing pressure of $P=150\,$bar in a nozzle with throat diameter $D^{*}=60\,\text{\textmu}$ µm. It is clear that the effective pumping speed needs to be larger than $R_{eff}>100\,$l/s to maintain the pressure in the differential chamber at the mbar level. When this condition is fulfilled, these estimations confirm that the equilibrium pressure in the main chamber remains at the level $P_{eq}^{(1)}\sim 10^{-5}\,$mbar, when the turbo-molecular pump is running at 4000 l/s. Here, we use the concept of effective pumping speed $R_{eff}$, where $R_{eff}$ is usually smaller than the nominal pumping speed $R_{nom}$ due to the finite conductance of the various elements of the pumping system (tubes, bellows, etc…) and also because the pumping speed varies with the operating pressure. This has to be carefully taken into account when designing the pumping system.
Once the differential pumping system had been set up, we measured the variation of the equilibrium pressure versus backing pressure for $\mathrm{N_{2}}$ and $\mathrm{H_{2}}$. Figure 5 shows the measurements: with the differential pumping apparatus, the pressure in the main chamber $P_{eq}^{(1)}$ remains below $3\times 10^{-4}\,$mbar, even with a backing pressure of 140140140\,140bar of $\mathrm{H_{2}}$ in the gas jet.
Figure 5: With differential pumping: equilibrium pressure in the main chamber versus backing pressure for a gas nozzle with throat diameter $D^{*}=60\,\text{\textmu}$ µm. The differential chamber is now pumped by a primary pump of 360360360\,360l/s.

### v3

@@ Abstract
Laser-Wakefield Accelerators (LWFA) running at kHz repetition rates hold great potential for applications. They typically operate with low-energy, highly compressed laser pulses focused in high-pressure gas targets. Experiments have shown that the best-quality electron beams are achieved using Hydrogen gas targets. However, continuous operation with Hydrogen requires a dedicated pumping system. In this work, we present a method for designing a differential pumping system, which we successfully implemented in our experiments. This enabled the first demonstration of continuous operation of a kHz LWFA using a high-pressure Hydrogen gas jet. The system effectively maintained a pressure below $3\times 10^{-4}\,$mbar, even with a free-flowing gas jet operating at 140 bar backing pressure. Numerical fluid dynamics and optical simulations were used to guide and validate the system’s design.
@@ I Introduction
In LWFA, the laser first propagates through a vacuum before being focused into a plasma, where the acceleration process occurs. The plasma is generated as the laser ionizes the gas supplied by the target system. Various types of gas targets are used in LWFA. Each type is designed to balance the required plasma density while minimizing gas leakage, which increases the pressure in the vacuum chamber where the laser propagates. Here are the most commonly used target systems: (i) gas jets: a supersonic gas flow expands from a nozzle into the vacuum chamber Semushin and Malka (2001); Schmid and Veisz (2012), (ii) gas cells: the gas is enclosed in a cell, creating an interaction volume in which the laser propagates through two apertures at the cell’s entrance and exit Osterhoff et al. (2008). These apertures must be small enough to minimize gas leakage in the main chamber, but large enough to avoid material ablation caused by the laser which would in turn negatively affect the accelerator performance. In order to mitigate the leaks, the gas cells operate at lower gas pressures, resulting in lower plasma densities compared to gas jets. Additionally, some gas cells utilize differential pumping, where the gas is removed from the cell simultaneously as it is introduced Kirchen et al. (2021); Drobniak et al. (2023). A last commonly used target system is the capillary discharge target, where the plasma is generated by a high-voltage discharge within a capillary filled with gas. There, the laser propagates through the guided mode of the plasma channel formed during plasma expansion and interaction with the capillary walls Leemans et al. (2006). This approach is relevant when long acceleration distances are needed (in the range of 111\,1cm), but the entrance capillary aperture can be damaged by high-intensity lasers. It is worth noticing that similar target systems are also being employed in the field of nuclear physics. Notably, some of their experiments make use of gas jets combined with differential pumping techniques, though these systems generally operate at much lower densities than those in the present study Ulbricht, Clausnitzer, and Graw (1972); Bittner, Kretschmer, and Schuster (1979); Treichel, Isenbügel, and Marquardt (1983); Shapira et al. (1985). Early-stage designs and implementations of such differential pumping systems have been explored in the field of high harmonic generation Sayrac et al. (2015) or LWFA Bohlen et al. (2022).
Advances in laser technology, along with the development of the target systems, have led to tremendous progress in the laser-plasma accelerator performances. Today, LWFA driven by 100 TW to PW scale laser systems can produce electron beams with energies ranging from 100100100\,100MeV to a few GeV over a few centimeters Mangles et al. (2004); Geddes et al. (2004); Faure et al. (2004); Kim et al. (2017); Oubrerie et al. (2022a); Miao et al. (2022). To date, the highest energy ever reached is 8 GeV using plasma discharges technology Gonsalves et al. (2019).
@@
Regarding vacuum conditions, utilizing a nitrogen plasma is convenient: each $\mathrm{N_{2}}$ molecule provides 10 electrons, enabling a backing pressure of only 20 to 30 bar in the gas jets to achieve the required plasma density. Additionally, heavy gases like nitrogen are easier to pump than light gases Gaede (1915); Becker (1966). However, ionization effects can significantly distort the laser pulse in $\mathrm{N_{2}}$ as different levels are ionizes at different positions within the laser pulse, causing a decrease in accelerator performance. To counter this, utilizing a light gas such as hydrogen is necessary Monzac et al. (2024). So far, kHz-scale experiments using an $\mathrm{H_{2}}$ plasma have been limited to burst mode operation as pumping $\mathrm{H_{2}}$ poses significant challenges Salehi et al. (2017, 2021). When a free-flowing gas jet of $\mathrm{H_{2}}$ is used at high pressure, it results in considerable gas loading in the vacuum chamber where the experiment is conducted. This phenomenon was observed by Salehi et al. Salehi et al. (2021), who reported a rapid pressure increase in the main chamber, rising from $2.6\times 10^{-2}\,$mbar to $2.0\times 10^{-1}\,$mbar in just 1 second when using a free-flowing gas jet with a $\sim 150\,\text{\textmu}$ µm FWHM aperture and a backing pressure of $\sim 30\,$bar. They operated their accelerator in burst mode to avoid gas loading in their experiments.
@@ II Differential pumping with gas jets and high pressure light gases
The optimal plasma density and length for a laser-plasma accelerator depend on the physical properties of the laser pulse. In the case of a few mJ, few fs pulse, the plasma density should typically be $n_{e}\sim 10^{19}-10^{20}\,\mathrm{cm}^{-3}$ and the plasma length $L=100-200\,\text{\textmu}$ µm Lu et al. (2006); Faure et al. (2018). To reach such plasma parameters, we use supersonic cylindrical gas nozzles and backing pressures ranging from a few tens of bar to about 150150150\,150bar. A cross-section diagram of the gas jet is shown in Figure 1b. In the following, we will refer to the gas nozzle parameters as follows: the diameter of the nozzle throat and nozzle opening are respectively written $D^{*}$ and $D$ and the corresponding throat and opening surfaces are $A^{*}$ and $A$, where we simply have $A=\pi D^{2}/4$. For example, a 60/180 gas nozzle has $D^{*}=60\,\text{\textmu}$ µm, $D=180\,\text{\textmu}$ µm, and provides the optimal plasma length and molecular density $n_{N_{2}}\sim 10^{19}\,\mathrm{cm}^{-3}$ when using $\mathrm{N_{2}}$ gas with pressure $P_{back}=30\,$bar backing pressure (see Figure 1a). The laser pulse is intense enough to ionize Nitrogen atoms to N5+, so that each $\mathrm{N_{2}}$ molecule releases 10 electrons, leading to the required plasma electron density of $n_{e}\sim 10^{20}\,\mathrm{cm}^{-3}$.
Figure 1: a) Electronic density $n_{e}$ of the plasma 150150150\,150µm above the gas jet as a function of the backing pressure in the gas jet and the nature of the gas for a 60/180 supersonic nozzle. b) Cross section of the supersonic gas jet, $D^{*}$ and $D$ are respectively the diameter of the nozzle throat and nozzle opening.
@@ II.1 Differential pumping use in high repetition rate LWFA
The pumping system is designed so that the residual equilibrium pressure in the experimental chamber, $P_{eq}^{(1)}$ stays sufficiently low and does not affect laser propagation. We set this limit empirically to $P_{eq}^{(1)}<10^{-2}\,$mbar since above this value, turbo-molecular pumps become inoperative, overheat, and the pumping system diverges. The nominal pumping speed of the pump, $R_{nom}$, is the theoretical pumping speed advertised by the manufacturer for a perfect pumping system. The effective pumping speed, $R_{eff}$ ($R_{eff}<R_{nom}$), is the real pumping speed in the system. The effective pumping speed is lower than the nominal speed primarily due to the presence of various components between the pump and the jet (belows and valves), which reduce the overall conductance of the pumping system. Additionally, leaks within the system contribute to a reduction in pumping efficiency, and the pumping speed varies with the operating pressure. The residual pressure at equilibrium $P_{eq}^{(1)}$ in the chamber can be estimated knowing the pumping speed $R_{eff}[\mathrm{l/s}]$ and the gas leak rate due to the gas jet $Q[\mathrm{mbar\cdot l/s}]$ by simply writing that:
@@
This allows to easily estimate the gas leak rate in our experiment. Typically, in Nitrogen, using a nozzle with $D^{*}=60\,\text{\textmu}$ µm and a backing pressure of $P_{back}=20\,$bar, we measure $P_{eq}^{(1)}=2.2\times 10^{-3}\,$mbar. Knowing that the pumping speed of our turbo-molecular pump is $R_{eff}=4000\,$l/s for Nitrogen at this pressure, we obtain that the gas leak is $Q=8.8\,\mathrm{mbar\cdot l/s}$.
Let us now calculate what happens if one uses $\mathrm{H_{2}}$ instead of Nitrogen. We assume complete ionization: each $\mathrm{H_{2}}$ molecule now releases two electrons instead of ten in $\mathrm{N_{2}}$. In addition, we use a 1D isentropic expansion model for compressible gas flows to extrapolate the molecular density in $\mathrm{H_{2}}$. The molecular density in the flow, $n_{mol}$, can be expressed according to the Boltzmann constant $k_{B}$, the Mach number $M_{Mach}$ and the initial pressure and temperature, $P_{0},\,T_{0}$ in the reservoir Zucker and Biblarz (2002):
@@
$Q_{\mathrm{H_{2},\,P_{0}=100bar}}=5Q_{\mathrm{N_{2},\,P_{0}=20bar}}\left(\frac{M_{N_{2}}}{M_{H_{2}}}\right)^{1/2}$ (3)
so that instead of $Q=8.8\,\mathrm{mbar\cdot l/s}$ for 20 bar of $\mathrm{N_{2}}$, the leak rate would be $Q=165\,\mathrm{mbar\cdot l/s}$ for 100 bar of $\mathrm{H_{2}}$. To compensate for that increase and keep the background pressure below $P_{eq}^{(1)}=3\times 10^{-3}\,$mbar, as when using $\mathrm{N_{2}}$, the pumping speed should be higher than $R_{eff}=5.5\times 10^{4}\,$l/s which is unrealistic as the largest turbo-molecular pumps on the market have maximum pumping speeds of 4000 l/s. Moreover, one should notice that $\mathrm{H_{2}}$ is more difficult to pump than $\mathrm{N_{2}}$. Overall, heavy gases are easier to pump than light gases; the compression ratio of the turbo-molecular pump goes as $K\propto\exp(\sqrt{M})$ with M the molar mass of the gas, and the pumping speed goes as $R\propto-\ln(K)$ Gaede (1915); Becker (1966), so the lighter the gas, the slower the pumping speed of the turbomolecular pump. This has been the bottleneck for using continuous flows of $\mathrm{He}$ or $\mathrm{H_{2}}$ in LWFA. For example, in our system, the pumping speed of our turbo-molecular pump drops down from 400040004000\,4000l/s in $\mathrm{N_{2}}$ to 300300300\,300l/s in $\mathrm{H_{2}}$ at $P=10^{-1}\,$mbar. In practice, this means that for $P_{0}=100\,$bar of $\mathrm{H_{2}}$, the equilibrium pressure would be around $P_{eq}^{(1)}\simeq 0.5\,$mbar, which is likely too high for the turbo-molecular pump to operate. Similar conclusions could be reached for $\mathrm{He}$.
Figure 2: Without differential pumping: measured equilibrium pressure in the main chamber versus backing pressure for a gas nozzle with nominal throat diameter $D^{*}=60\,\text{\textmu}$ µm (measured throat diameter: $D^{*}=40\pm 2\,\text{\textmu}$ µm).
This has been verified experimentally: we measured the pressure in the vacuum chamber in which a supersonic nozzle lets out a continuous flow of gas. The vacuum chamber is pumped by a $R_{eff}=4000\,$l/s turbo-molecular pump connected to a primary pump with pumping speed $R_{nom}=130\,$l/s. In Figure 2, we show the variation of the equilibrium pressure versus backing pressure for $\mathrm{N_{2}}$, and $\mathrm{He}$. The results show very clearly that for $\mathrm{He}$, the equilibrium pressure increases abruptly above $P_{back}=20\,$bar and the pumping system diverges above that, thus precluding operation at high $\mathrm{He}$ flow. This clearly confirms that a specific solution should be implemented for pumping high-pressure ($P>100\,$bar) light gases.
@@ II.2 Design of the differential pumping system
An alternative solution is to use a differential pumping system: the gas jet is enclosed into a small vacuum chamber, the differential chamber, that is pumped by a powerful primary pump able to work efficiently at the mbar level with a pumping speed of $R_{nom}=360\,$l/s. The idea is to pump out most of the gas from the small chamber while tolerating a higher equilibrium pressure at the mbar level, $P_{eq}^{(2)}\sim\,$mbar within the differential chamber subsystem. Then, by restricting the leakage from the differential chamber to the main chamber, the pressure can be kept below $P_{eq}^{(1)}<10^{-3}\,$mbar in the main chamber. In this case, the laser propagation is not affected by the residual gas, provided that the mbar region in the differential chamber is sufficiently short, for example in our case millimeter scale. Figure 3 shows a conceptual schematic of the differential scheme that is implemented in the experiment. The laser enters and exits the chamber via two cones that terminate with 111\,1mm holes for restricting gas leakage. In addition, the “high” density region is limited to a $4$-mm gap in between the two cones to limit its impact on laser propagation.
@@
The experimental measurement of the gas leak rate $Q$ can be compared to a theoretical estimation using well-known fluid gas dynamics Zucker and Biblarz (2002). An estimate of the mass flow rate can be obtained using a 1D isentropic model of the gas flow in the nozzle. For a 1D stationary and turbulent flow, the mass flow rate $\dot{m}=\rho\nu A$ is constant throughout the flow; at the nozzle throat, it reduces to:
@@
where $\rho^{*},\,\nu^{*}$ and $A^{*}$ are respectively the mass density, fluid velocity, and the surface at the nozzle throat, $P_{0}$ and $T_{0}$ are the pressure and temperature in the reservoir. Note that here, the 0 subscript refers to the reservoir, i.e. the stagnation volume where the fluid velocity is $\nu_{0}=0$. The leak rate in $\mathrm{mbar\cdot l/s}$ can then be obtained from $Q\,=PdV/dt=k_{B}T\frac{dN}{dt}$ which can thus be written as:
@@
with $R=R_{0}/M$, $R_{0}=8.314\,$J/mol/K is the ideal gas constant and $M$ is the molar mass of the fluid. Taking $P_{back}=P_{0}=20\,$bar and $T=T_{0}=300\,$K for a nozzle with $D^{*}=60\,\text{\textmu}$ µm, we obtain $Q=11.6\,\mathrm{mbar\cdot l/s}$. The nominal throat diameter of the jets is $D^{*}=60\,\text{\textmu}$ µm. However, the actual throat diameter deviates from the generic requested value; for the dataset shown in Figure 2 the measured throat diameter is $D^{*}=40\pm 2\,\text{\textmu}$ µm, leading to $Q=5.1\,\mathrm{mbar\cdot l/s}$ which is close to the experimental value found earlier ($Q_{exp}=8.8\,\mathrm{mbar\cdot l/s}$). This analysis validates the fact that this simple model can be used for designing the pumping system as it provides a pressure estimate that matches the experimental measurement within a factor of two.
@@
Using this model, it is straightforward to determine $P_{eq}^{(1)}$ and $P_{eq}^{(2)}$, respectively the equilibrium pressures in the experimental and differential chambers, that are reached as a function of the pumping speed $R_{eff}$ used for pumping the differential chamber. Figure 4 shows the results for a backing pressure of $P=150\,$bar in a nozzle with throat diameter $D^{*}=60\,\text{\textmu}$ µm. It is clear that the effective pumping speed needs to be larger than $R_{eff}>100\,$l/s to maintain the pressure in the differential chamber at the mbar level. When this condition is fulfilled, these estimates confirm that the equilibrium pressure in the main chamber remains at the level $P_{eq}^{(1)}\sim 10^{-5}\,$mbar, where the turbo-molecular pump works in optimal conditions.
Once the differential pumping system was set up, we measured the variation of the equilibrium pressure versus backing pressure for $\mathrm{N_{2}}$ and $\mathrm{H_{2}}$. Figure 5 shows the pressure measurements in the main chamber with the differential pumping apparatus in operation. The observed discrepancy between the experimental measurements and the theoretical predictions shown in Figure 4 arises because the mathematical model inherently simplifies or neglects certain real-world factors that influence the system’s performance, such as leaks in the chamber or deviations of the actual conductance from a geometrically simplified model. Experimentally, the lowest pressure achieved in the main chamber without load is approximately $5\times 10^{-5}\,$mbar, which is significantly higher than the model’s prediction of $2.5\times 10^{-6}\,$mbar. Nevertheless, the equilibrium pressure in the main chamber $P_{eq}^{(1)}$ remains below $3\times 10^{-4}\,$mbar, even with a backing pressure of 140140140\,140bar of $\mathrm{H_{2}}$ in the gas jet.
Figure 5: With differential pumping: equilibrium pressure in the main chamber versus backing pressure for a gas nozzle with throat diameter $D^{*}=60\,\text{\textmu}$ µm. The differential chamber is pumped by a primary pump of 360360360\,360l/s.